# Investigating the Event-Shape Methods in Search for the Chiral Magnetic Effect in Relativistic Heavy Ion Collisions

### v3

@@ IV.3 Toy Model Verification
Figure A.14 shows the ESE results from Toy Model I (default 2% $K_{S}$), Toy Model I (10% $K_{S}$), and Toy Model II (mass continuum) with POIs from $10.3<|\eta|<$ 1 and $q_{2}^{2}$ from $|\eta|<0.3$. There seems to be an indication of nonlinear behavior in the Toy Model II result, similar to that in the ESS result in Fig. A.14 right panel.
Figure 8: Toy model ESS and ESE intercepts. Shown are intercepts (upper panel) and intercepts divided by the inclusive $\Delta\gamma$ value (lower panel) from the ESS (blue points) and ESE (red and pink points) methods using the toy models. Toy model inputs are from parameterizations to experimental data of 30–40% Au+Au collisions at 200 GeV. For ESS, the POI acceptance is $|\eta|<1$ and $20.2<p_{\rm T}<$ 2 GeV/$c$, and the $\hat{q}_{2,{\rm pair}}^{2}$ is always calculated from the same POIs. For ESE, the POI acceptance is $10.3<|\eta|<$ 1, and the $\hat{q}_{2}^{2}$ is calculated from particles in $|\eta|<0.3$ (red points) or $43<|\eta|<$ 4 for three of the studied cases (pink points); The $p_{\rm T}$ range for both POIs and particles used for $\hat{q}_{2}^{2}$ calculations is $20.2<p_{\rm T}<$ 2 GeV/$c$. The rightmost red point is outside the frame.
@@
The intercepts from the ESE method are shown in the red markers. The statistical precision is high because the toy models are computationally inexpensive. The ESE intercepts are consistent with zero for the two lower $K_{S}$-contamination cases studied in Toy Model I. However, the ESE intercepts from the two higher $K_{S}$-contamination cases of Toy Model I (positive) and the intercept from Toy Model II (negative) are inconsistent with zero. We postulate the cause to be nonflow correlations between POIs in $10.3<|\eta|<$ 1 and the $q_{2}^{2}$ calculated from $|\eta|<0.3$ in our toy model studies because resonance decay daughters can end up in both regions which are not far apart. (We have mentioned this nonflow effect in the $q_{2}^{2}$ event selection in our physics model discussion of Fig. 5, though the statistical uncertainties there are large.) In addition, the effect of adding a mass continuum is different from simply increasing the $K_{S}$ contamination, presumably because of the differing decay kinematics and opening angles.
To reduce this nonflow effect, we classify events in the ESE method using $q_{2}^{2}$ calculated from forward/backward pseudorapidity regions of $43<|\eta|<$ 4, far displaced from the POI’s region of $10.3<|\eta|<$ 1. The results are shown in Fig. A.14 for Toy Model I (20% $K_{S}$), Toy Model I (30% $K_{S}$), and Toy Model II (mass continuum). The ESE intercepts and the intercepts divided by the overall $\Delta\gamma$ values are shown in the pink markers in Fig. 8 with good statistical precision. These intercepts are consistent with zero, as one would expect in the absence of nonflow. These results highlight the importance to use $q_{2}^{2}$ calculated in pseudorapidity range far distanced from that of the POIs.
@@
The charge-dependent azimuthal correlator $\Delta\gamma$ is commonly used to search for the chiral magnetic effect (CME), a parity and charge-parity violating phenomenon predicted by quantum chromodynamics. The $\Delta\gamma$ correlator is contaminated by a major physics background induced by elliptic flow anisotropy $v_{2}$ in relativistic heavy ion collisions. The event-shape engineering (ESE) and event-shape selection (ESS) methods have been exploited to search for the CME by projecting $\Delta\gamma$ to the measured $v_{2}$ of vanishing value, relying on dynamical fluctuations and statistical fluctuations of $v_{2}$, respectively.
@@
It is further found, using toy models, that $\Delta\gamma_{\textsc{ess}}$ appears to depend on the details of the event content, for example, the relative abundances of resonances and the presence of a mass continuum. Given that the event-shape selection variable $\hat{q}_{2,{\rm pair}}^{2}$, the single particle $v_{2,{\rm single}}$, and the three-point correlator $\Delta\gamma$ in $\hat{q}_{2,{\rm pair}}^{2}$-classified events all depend on the particles of interest (POIs), which are made of not only primordial particles but also all decay products of electromagnetic- and strong-decay resonances and partial products of weak-decay resonances, we conjecture that the ESS intercept $\Delta\gamma_{\textsc{ess}}$ is a complicated function of all elements that make up the event. These elements include resonance abundances relative to primordial pions, resonance $v_{2}$’s and $p_{\rm T}$ spectra, as well as primordial particle $v_{2}$ and $p_{\rm T}$ spectra. One may postulate a diagram such as in Fig. 9 where $\Delta\gamma_{\textsc{ess}}$ is depicted as a function of, for the sake of visualization, two elements of the event. For some combinations of the elements, $\Delta\gamma_{\textsc{ess}}$ is negative and for some others, $\Delta\gamma_{\textsc{ess}}$ is positive; for some accidental combinations, $\Delta\gamma_{\textsc{ess}}$ can happen to be zero. The issue is, unfortunately, that it is unknown, needless to say proven, what combinations of all those elements in the event would give the ideal $\Delta\gamma_{\textsc{ess}}=0$ in the absence of CME. This makes it impractical to search for CME using the ESS method.
Figure 9: Postulated behavior (for illustration only) of the ESS intercept $\Delta\gamma_{\textsc{ess}}$ on event elements of many dimensions, represented here for visualization purpose only by two, labeled as “resonance 1” and “resonance 2.” In the absence of CME, the value of $\Delta\gamma_{\textsc{ess}}$ can be positive or negative, and for some combinations of elements, can happen to be zero. This is because of the intertwining $\Delta\gamma$, $v_{2,{\rm single}}$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}$, all of which are calculated from the same set of POIs. This makes the ESS method highly model-dependent and unpractical for rigorous CME searches.
@@ Appendix A Simulation Plots
Figure A.1: avfd ESS results. Three values of $n_{5}/s=0$ (left panel), 0.1 (center panel), and 0.2 (right panel) are implemented in the avfd simulations of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV, with approximately $2\times 10^{7}$ events each for the 30–40% centrality. The $\Delta\gamma$ is plotted as a function of $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ (Eqs. 5,9). POIs are from acceptance $|\eta|<1$ and $20.2<p_{\rm T}<$ 2 GeV/$c$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ is computed from the same POIs. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $v_{2,{\rm single}}$ (Eq. 10). The red line is a first-order polynomial fit in the range of $0.20<v_{2,{\rm single}}<$ 0.2.
@@
Figure A.2: avfd ESE results. The same avfd data as in Fig. A.2 is used. The $\Delta\gamma$ is plotted as a function of $mean{v_{2}}$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $20.3<|\eta|<$ 2, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $20.2<p_{\rm T}<$ 2 GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10). The red line is a first-order polynomial fit to all data points.
@@
Figure A.3: ampt ESS results. Shown are three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV (upper panels) and at 27 GeV (lower panels) simulated by ampt, with approximately $4.2\times 10^{8}$ and $1.8\times 10^{9}$ events for each centrality at each energy. The $\Delta\gamma$ is plotted as a function of $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ (Eqs. 5,9). POIs are from acceptance $|\eta|<1$ and $20.2<p_{\rm T}<$ 2 GeV/$c$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ is computed from the same POIs. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $v_{2,{\rm single}}$ (Eq. 10). The red line is a first-order polynomial fit in the range of $0.20<v_{2,{\rm single}}<$ 0.2
@@
Figure A.4: ampt ESE results. The same ampt data as in Fig. A.3 is used. The $\Delta\gamma$ is plotted as a function of $\left\langle v_{2}\right\rangle$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $10.3<|\eta|<$ 1, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $20.2<p_{\rm T}<$ 2 GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10). The red line is a first-order polynomial fit to all data points.
@@
Figure A.5: ampt ESE results with forward/backward $\hat{q}_{2}^{2}\{2\}$. As same as Fig. A.5 except that the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in the forward/backward pseudorapidity region of $53<|\eta|<$ 5 for 200 GeV and $32<|\eta|<$ 3 for 27 GeV.
@@
Figure A.6: epos4 ESS results. Shown are three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV (upper panels) and at 27 GeV (lower panels) simulated by epos4, with approximately $1.6\times 10^{7}$ and $8.5\times 10^{6}$ events for each centrality, respectively. The $\Delta\gamma$ is plotted as a function of $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ (Eqs. 5,9). POIs are from acceptance $|\eta|<2$ and $20.2<p_{\rm T}<$ 2 GeV/$c$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ is computed from the same POIs. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $v_{2,{\rm single}}$ (Eq. 10).
@@
Figure A.7: epos4 ESE results. The same epos4 data as in Fig. A.7 is used. The $\Delta\gamma$ is plotted as a function of $\left\langle v_{2}\right\rangle$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $20.3<|\eta|<$ 2, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $20.2<p_{\rm T}<$ 2 GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10).
@@
Figure A.8: hydjet++ ESS results. Shown are three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV (upper panels) and at 27 GeV (lower panels) simulated by hydjet++, with approximately $2.5\times 10^{8}$ and $5.7\times 10^{8}$ events for each centrality, respectively. The $\Delta\gamma$ is plotted as a function of $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ (Eqs. 5,9). POIs are from acceptance $|\eta|<1$ and $20.2<p_{\rm T}<$ 2 GeV/$c$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ is computed from the same POIs. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $v_{2,{\rm single}}$ (Eq. 10).
@@
Figure A.9: hydjet++ ESE results. The same hydjet++ data as in Fig. A.9 is used. The $\Delta\gamma$ is plotted as a function of $\left\langle v_{2}\right\rangle$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $10.3<|\eta|<$ 1, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $20.2<p_{\rm T}<$ 2 GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10).
@@
Figure A.11: hydjet++ ESE results (forward/backward rapidity $\hat{q}_{2}^{2}$). As same as Fig. A.9 except that the event selection variable $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6) is computed from particles in forward/backward pseudorapidity regions of $53<|\eta|<$ 5 for 200 GeV and $32<|\eta|<$ 3 for 27 GeV.
Figure A.12: Toy model ESS results. Toy Model I with the accepted $K_{S}$ fraction of the default 2% (left panel) and of 10% (middle panel), and Toy Model II with mass continuum (right panel). The $K_{S}$, $\rho^{0}$, and primordial pion inputs are from parameterizations to experimental data of 30–40% centrality Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV. The numbers of events are $2\times 10^{8}$ for each panel. The $\Delta\gamma$ is plotted as a function of $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ (Eqs. 5,9). POIs are from acceptance $|\eta|<1$ and $20.2<p_{\rm T}<$ 2 GeV/$c$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ is computed from the same POIs. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $v_{2,{\rm single}}$ (Eq. 10). The red line is a first-order polynomial fit in the range of $0.20<v_{2,{\rm single}}<$ 0.2.
Figure A.13: Toy model ESE results. Toy Model I with the accepted $K_{S}$ fraction of the default 2% (left panel, $2.8\times 10^{9}$ events) and of 10% (middle panel, $2.0\times 10^{9}$ events), and Toy Model II with mass continuum (right panel, $2.2\times 10^{8}$ events). The $\Delta\gamma$ is plotted as a function of $\left\langle v_{2}\right\rangle$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $10.3<|\eta|<$ 1, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $20.2<p_{\rm T}<$ 2 GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10). Other information is as same as those in Fig. A.14. The red line is a first-order polynomial fit to all data points.
Figure A.14: Toy model ESE results (forward/backward rapidity $\hat{q}_{2}^{2}$). Toy Model I with the accepted $K_{S}$ fraction of 20% (left panel, $2.2\times 10^{8}$ events) and of 30% (middle panel, $2.2\times 10^{8}$ events), and Toy Model II with mass continuum (right panel, $10^{8}$ events). The event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $43<|\eta|<$ 4. Other information is as same as those in Fig. A.14.

### v4

@@ IV.3 Toy Model Verification
Figure A.14 shows the ESE results from Toy Model I (default 2% $K_{S}$), Toy Model I (10% $K_{S}$), and Toy Model II (mass continuum) with POIs from $0.3<|\eta|<1$ and $q_{2}^{2}$ from $|\eta|<0.3$. There seems to be an indication of nonlinear behavior in the Toy Model II result, similar to that in the ESS result in Fig. A.14 right panel.
Figure 8: Toy model ESS and ESE intercepts. Shown are intercepts (upper panel) and intercepts divided by the inclusive $\Delta\gamma$ value (lower panel) from the ESS (blue points) and ESE (red and pink points) methods using the toy models. Toy model inputs are from parameterizations to experimental data of 30–40% Au+Au collisions at 200 GeV. For ESS, the POI acceptance is $|\eta|<1$ and $0.2<p_{\rm T}<2$ GeV/$c$, and the $\hat{q}_{2,{\rm pair}}^{2}$ is always calculated from the same POIs. For ESE, the POI acceptance is $0.3<|\eta|<1$, and the $\hat{q}_{2}^{2}$ is calculated from particles in $|\eta|<0.3$ (red points) or $3<|\eta|<4$ for three of the studied cases (pink points); The $p_{\rm T}$ range for both POIs and particles used for $\hat{q}_{2}^{2}$ calculations is $0.2<p_{\rm T}<2$ GeV/$c$. The rightmost red point is outside the frame.
@@
The intercepts from the ESE method are shown in the red markers. The statistical precision is high because the toy models are computationally inexpensive. The ESE intercepts are consistent with zero for the two lower $K_{S}$-contamination cases studied in Toy Model I. However, the ESE intercepts from the two higher $K_{S}$-contamination cases of Toy Model I (positive) and the intercept from Toy Model II (negative) are inconsistent with zero. We postulate the cause to be nonflow correlations between POIs in $0.3<|\eta|<1$ and the $q_{2}^{2}$ calculated from $|\eta|<0.3$ in our toy model studies because resonance decay daughters can end up in both regions which are not far apart. (We have mentioned this nonflow effect in the $q_{2}^{2}$ event selection in our physics model discussion of Fig. 5, though the statistical uncertainties there are large.) In addition, the effect of adding a mass continuum is different from simply increasing the $K_{S}$ contamination, presumably because of the differing decay kinematics and opening angles.
To reduce this nonflow effect, we classify events in the ESE method using $q_{2}^{2}$ calculated from forward/backward pseudorapidity regions of $3<|\eta|<4$, far displaced from the POI’s region of $0.3<|\eta|<1$. The results are shown in Fig. A.14 for Toy Model I (20% $K_{S}$), Toy Model I (30% $K_{S}$), and Toy Model II (mass continuum). The ESE intercepts and the intercepts divided by the overall $\Delta\gamma$ values are shown in the pink markers in Fig. 8 with good statistical precision. These intercepts are consistent with zero, as one would expect in the absence of nonflow. These results highlight the importance to use $q_{2}^{2}$ calculated in pseudorapidity range far distanced from that of the POIs.
@@
The charge-dependent azimuthal correlator $\Delta\gamma$ is commonly used to search for the chiral magnetic effect (CME), a parity and charge-parity violating phenomenon predicted by quantum chromodynamics. The $\Delta\gamma$ correlator is affected by a major physics background induced by elliptic flow anisotropy $v_{2}$ in relativistic heavy ion collisions. The event-shape engineering (ESE) and event-shape selection (ESS) methods have been exploited to search for the CME by projecting $\Delta\gamma$ to the measured $v_{2}$ of vanishing value, relying on dynamical fluctuations and statistical fluctuations of $v_{2}$, respectively.
@@
It is further found, using toy models, that $\Delta\gamma_{\textsc{ess}}$ appears to depend on the details of the event content, for example, the relative abundances of resonances and the presence of a mass continuum. Given that the event-shape selection variable $\hat{q}_{2,{\rm pair}}^{2}$, the single particle $v_{2,{\rm single}}$, and the three-point correlator $\Delta\gamma$ in $\hat{q}_{2,{\rm pair}}^{2}$-classified events all depend on the particles of interest (POIs), which are made of not only primordial particles but also all decay products of electromagnetic- and strong-decay resonances and partial products of weak-decay resonances, we conjecture that the ESS intercept $\Delta\gamma_{\textsc{ess}}$ is a complicated function of all elements that make up the event. These elements include resonance abundances relative to primordial pions, resonance $v_{2}$’s and $p_{\rm T}$ spectra, as well as primordial particle $v_{2}$ and $p_{\rm T}$ spectra. Depending on the detailed physics of these elements, $\Delta\gamma_{\textsc{ess}}$ can be positive, negative, or zero. The issue is, unfortunately, that it is unknown, needless to say proven, what combinations of all those elements in the event would give the ideal $\Delta\gamma_{\textsc{ess}}=0$ in the absence of CME. This makes it impractical to search for CME using the ESS method.
@@ Appendix A Simulation Plots
Figure A.1: avfd ESS results. Three values of $n_{5}/s=0$ (left panel), 0.1 (center panel), and 0.2 (right panel) are implemented in the avfd simulations of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV, with approximately $2\times 10^{7}$ events each for the 30–40% centrality. The $\Delta\gamma$ is plotted as a function of $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ (Eqs. 5,9). POIs are from acceptance $|\eta|<1$ and $0.2<p_{\rm T}<2$ GeV/$c$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ is computed from the same POIs. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $v_{2,{\rm single}}$ (Eq. 10). The red line is a first-order polynomial fit in the range of $0<v_{2,{\rm single}}<0.2$.
@@
Figure A.2: avfd ESE results. The same avfd data as in Fig. A.2 is used. The $\Delta\gamma$ is plotted as a function of $mean{v_{2}}$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $0.3<|\eta|<2$, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $0.2<p_{\rm T}<2$ GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10). The red line is a first-order polynomial fit to all data points.
@@
Figure A.3: ampt ESS results. Shown are three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV (upper panels) and at 27 GeV (lower panels) simulated by ampt, with approximately $4.2\times 10^{8}$ and $1.8\times 10^{9}$ events for each centrality at each energy. The $\Delta\gamma$ is plotted as a function of $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ (Eqs. 5,9). POIs are from acceptance $|\eta|<1$ and $0.2<p_{\rm T}<2$ GeV/$c$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ is computed from the same POIs. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $v_{2,{\rm single}}$ (Eq. 10). The red line is a first-order polynomial fit in the range of $0<v_{2,{\rm single}}<0.2$
@@
Figure A.4: ampt ESE results. The same ampt data as in Fig. A.3 is used. The $\Delta\gamma$ is plotted as a function of $\left\langle v_{2}\right\rangle$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $0.3<|\eta|<1$, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $0.2<p_{\rm T}<2$ GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10). The red line is a first-order polynomial fit to all data points.
@@
Figure A.5: ampt ESE results with forward/backward $\hat{q}_{2}^{2}\{2\}$. As same as Fig. A.5 except that the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in the forward/backward pseudorapidity region of $3<|\eta|<5$ for 200 GeV and $2<|\eta|<3$ for 27 GeV.
@@
Figure A.6: epos4 ESS results. Shown are three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV (upper panels) and at 27 GeV (lower panels) simulated by epos4, with approximately $1.6\times 10^{7}$ and $8.5\times 10^{6}$ events for each centrality, respectively. The $\Delta\gamma$ is plotted as a function of $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ (Eqs. 5,9). POIs are from acceptance $|\eta|<2$ and $0.2<p_{\rm T}<2$ GeV/$c$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ is computed from the same POIs. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $v_{2,{\rm single}}$ (Eq. 10).
@@
Figure A.7: epos4 ESE results. The same epos4 data as in Fig. A.7 is used. The $\Delta\gamma$ is plotted as a function of $\left\langle v_{2}\right\rangle$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $0.3<|\eta|<2$, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $0.2<p_{\rm T}<2$ GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10).
@@
Figure A.8: hydjet++ ESS results. Shown are three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV (upper panels) and at 27 GeV (lower panels) simulated by hydjet++, with approximately $2.5\times 10^{8}$ and $5.7\times 10^{8}$ events for each centrality, respectively. The $\Delta\gamma$ is plotted as a function of $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ (Eqs. 5,9). POIs are from acceptance $|\eta|<1$ and $0.2<p_{\rm T}<2$ GeV/$c$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ is computed from the same POIs. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $v_{2,{\rm single}}$ (Eq. 10).
@@
Figure A.9: hydjet++ ESE results. The same hydjet++ data as in Fig. A.9 is used. The $\Delta\gamma$ is plotted as a function of $\left\langle v_{2}\right\rangle$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $0.3<|\eta|<1$, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $0.2<p_{\rm T}<2$ GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10).
@@
Figure A.11: hydjet++ ESE results (forward/backward rapidity $\hat{q}_{2}^{2}$). As same as Fig. A.9 except that the event selection variable $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6) is computed from particles in forward/backward pseudorapidity regions of $3<|\eta|<5$ for 200 GeV and $2<|\eta|<3$ for 27 GeV.
Figure A.12: Toy model ESS results. Toy Model I with the accepted $K_{S}$ fraction of the default 2% (left panel) and of 10% (middle panel), and Toy Model II with mass continuum (right panel). The $K_{S}$, $\rho^{0}$, and primordial pion inputs are from parameterizations to experimental data of 30–40% centrality Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV. The numbers of events are $2\times 10^{8}$ for each panel. The $\Delta\gamma$ is plotted as a function of $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ (Eqs. 5,9). POIs are from acceptance $|\eta|<1$ and $0.2<p_{\rm T}<2$ GeV/$c$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ is computed from the same POIs. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $v_{2,{\rm single}}$ (Eq. 10). The red line is a first-order polynomial fit in the range of $0<v_{2,{\rm single}}<0.2$.
Figure A.13: Toy model ESE results. Toy Model I with the accepted $K_{S}$ fraction of the default 2% (left panel, $2.8\times 10^{9}$ events) and of 10% (middle panel, $2.0\times 10^{9}$ events), and Toy Model II with mass continuum (right panel, $2.2\times 10^{8}$ events). The $\Delta\gamma$ is plotted as a function of $\left\langle v_{2}\right\rangle$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $0.3<|\eta|<1$, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $0.2<p_{\rm T}<2$ GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10). Other information is as same as those in Fig. A.14. The red line is a first-order polynomial fit to all data points.
Figure A.14: Toy model ESE results (forward/backward rapidity $\hat{q}_{2}^{2}$). Toy Model I with the accepted $K_{S}$ fraction of 20% (left panel, $2.2\times 10^{8}$ events) and of 30% (middle panel, $2.2\times 10^{8}$ events), and Toy Model II with mass continuum (right panel, $10^{8}$ events). The event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $3<|\eta|<4$. Other information is as same as those in Fig. A.14.